OpenStudy (anonymous):

PLEASE HELP ME!!!! Solve: √(x-5)<7 I believe it's either: x<54 or 5≤x<54

jimthompson5910 (jim_thompson5910):

First we need to find the domain

jimthompson5910 (jim_thompson5910):

since you cannot take the square root of a negative number, this means x-5 can't be negative, so this means x - 5 >= 0 solve for x x - 5 >= 0 x-5+5 >= 0+5 x >= 5

jimthompson5910 (jim_thompson5910):

now let's solve for x in sqrt(x-5) < 7 sqrt(x-5) < 7 ( sqrt(x-5) )^2 < ( 7 )^2 ... square both sides | x - 5 | < 49 -49 < x - 5 < 49 -49+5 < x < 49+5 -44 < x < 54 5 <= x < 54 ... applying the domain restriction here

jimthompson5910 (jim_thompson5910):

in line 3, I used the idea that ( sqrt(x) )^2 = |x| and in line 4, I used the property that if |x| < k and k is some positive number, then -k < x < k

OpenStudy (anonymous):

okay thank you so much, so in final 5≤x<54, but if x is less than 5 the solution is not real but it would still work?

jimthompson5910 (jim_thompson5910):

if x was less than 5, then x-5 would be negative (try x = 4 and you'll get x-5 = 4-5 = -1) but again you cannot take the square root of a negative number

jimthompson5910 (jim_thompson5910):

that's why the domain of f(x) = sqrt(x-5) is x>= 5

OpenStudy (anonymous):

so it has to be greater than or equal to 5 for the solution to be a real number and less than 54 for the solution to be less than 7, so you can't include imaginary numbers in inequalities?

jimthompson5910 (jim_thompson5910):

you can't include imaginary numbers in inequalities because you can't naturally order imaginary or complex numbers ex: which number is bigger, 2+i or 3-i ? there's no way of sorting the two like you can sort real numbers

jimthompson5910 (jim_thompson5910):

so inequalities only apply to real numbers

OpenStudy (anonymous):

okay thank you so much!!!

jimthompson5910 (jim_thompson5910):

you're welcome